# One Backpropagation in Two Tower Recommendation Models

### v3

@@ 5.5. Exploration Study
We next explore the recommendation effects of learning diverse interests for users. Fig. 7 reports the Precision (P) and Recall (R) statistics in each latent type for the selected user in Fig. 4. The percentage means the ratio of items of the same type over the length of the recommendation list. The left is by using the traditional TwoBP, where three latent types are identified for his testing items; The right is by using our OneBP, where four latent types are identified. Comparing the Precision and Recall for majority testing items, our OneBP outperforms the TwoBP. For latent types of minority testing items, our OneBP also performs better, so achieving better overall recommendation results. Fig. 8 reports the top-$K$ recommendation statistics over all users. With the increase of $K$, our OneBP presents a clearer tendency of much reduced majority items. Furthermore, minority items are distributed more proportionally in different clusters. This suggests a better diversification for items. While such diversification gets paid in terms better recommendation performance for minority items in each latent type. As such, our OneBP achieves better overall performance than the TwoBP.

### v4

@@ 5.5. Exploration Study
We next explore the recommendation effects of learning diverse interests for users. Fig. 7 reports the Precision (P) and Recall (R) statistics in each latent type for the selected user in Fig. 4. The percentage means the ratio of items of the same type over the length of the recommendation list. The left is by using the traditional TwoBP, where three latent types are identified for his/her testing items; The right is by using our OneBP, where four latent types are identified. Comparing the Precision and Recall for majority testing items, our OneBP outperforms the TwoBP. For latent types of minority testing items, our OneBP also performs better, so achieving better overall recommendation results. Fig. 6 reports the top-$K$ recommendation statistics over all users. With the increase of $K$, our OneBP presents a clearer tendency of much reduced majority items. Furthermore, minority items are distributed more proportionally in different clusters. This suggests a better diversification for items. While such diversification gets paid in terms better recommendation performance for minority items in each latent type. As such, our OneBP achieves better overall performance than the TwoBP.